I have to do this with the substitution method: Help me please
3x = 4y
x = (2y+1)/3

I know the answer is x=2/3 and y = 1/2 but I don't know the method
Thank you

Answer by umnikmo(75)

from the first we can find x in terms of y: x=4y/3, now we substitute this into second equation, we get:
4y/3=(2y+1)/3
we can multiply both sides by 3:
4y=2y+1
2y=1
y=1/2.
now, as x=4y/3=4*(1/2)/3=2/3
same as your answers.